OpenStudy (anonymous):

Evaluate. C(9, 7) A. 181,440 B. 72 C. 36 D. 63

OpenStudy (damonte212):

B.

OpenStudy (anonymous):

Thanks

OpenStudy (damonte212):

np

OpenStudy (igreen):

@damonte212 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (igreen):

\(\sf C(m, n) = \dfrac{m!}{(n!)(m-n)!}\)

OpenStudy (igreen):

\(\sf C(9, 7) = \dfrac{9!}{(7!)(9-7)!}\) Simplify

OpenStudy (igreen):

Btw, B is incorrect @kyrsten6372

OpenStudy (igreen):

\(\sf C(9, 7) = \dfrac{9!}{(7!)(2)!}\)

OpenStudy (anonymous):

Okay

OpenStudy (igreen):

\(\sf C(9, 7) = \dfrac{9 \times 8 \times \cancel 7 \times \cancel 6 \times \cancel 5 \times \cancel 4 \times \cancel 3 \times \cancel 2}{\cancel 7 \times \cancel 6 \times \cancel 5 \times \cancel 4 \times \cancel 3 \times \cancel2 \times 2}\)

OpenStudy (igreen):

Can you finish it?

OpenStudy (anonymous):

72

OpenStudy (igreen):

The top equals 72, but there's still a 2 at the bottom.

OpenStudy (anonymous):

36

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

Thanks